# scale cd/.style=every label/.append style=scale=#1, cells=nodes=scale=#1

### v1

@@ 8. Brauer pairs for splendid Rickard complexes
where for $\nu=(\sigma,(S,a),\tau))$, we set $X(\nu):=N_{I\times H}(\Delta(P,\sigma,S),e\otimes a^{*})$, $Y(\nu):=N_{H\times K}(\Delta(S,\tau,R),a\otimes d^{*})$, $C(\nu):=eC(\Delta(P,\sigma,S))a$ and $C^{*}(\nu):=aC^{*}(\Delta(S,\tau,R))d$. In particular, when $\nu=(\phi,(Q,f),c_{t}\circ\psi^{-1})$, we have $X(\nu)=X$, $Y(\nu)={}^{(t,1)}Y$, $C(\nu)=C(\lambda)$ and, by Proposition 3.4, $C^{*}(\nu)\cong{}^{(t,1)}(C(\omega)^{*})$. Thus, since $(\phi,(Q,f),c_{t}\circ\psi^{-1})\in\hat{\Omega}$, we conclude that
$\operatorname{Ind}_{X\ast{}^{(t,1)}Y}^{N_{I\times K}(\Delta(P,\phi\circ c_{t}\circ\psi^{-1},R))}\left(C(\lambda)\mathop{\otimes}_{kH}^{X,{}^{(t,1)}Y}{}^{(t,1)}(C(\omega)^{*})\right)$
@@
As a corollary of the preceding theorem, we obtain that the Brauer pairs for a splendid Rickard equivalence and the Brauer pairs for its Lefschetz invariant coincide.
@@
Let $C$ be a splendid Rickard complex between $A$ and $B$ and let $\gamma:=\Lambda(C)$ be its corresponding $p$-permutation equivalence. Then $\mathcal{BP}(C)=\mathcal{BP}(\gamma)$. In particular,
@@
By Proposition 7.7 and Theorem 8.4, we may assume that $A$ and $B$ are blocks. Then the fact that $\mathcal{BP}(C)\subseteq\mathcal{BP}(\gamma)$ follows from Theorem 8.7. The other inclusion $\mathcal{BP}(\gamma)\subseteq\mathcal{BP}(C)$ was already established at the beginning of this section. Claims (a)-(c) above follow immediately from [4, Theorem 10.11]. Claim (d) follows from [4, Theorem 11.2]. ∎
Another easy corollary we obtain is that splendid Rickard complexes induce local normalizer splendid equivalences.
@@
Let $C$ be a splendid Rickard complex between blocks $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $C$-Brauer pair. Set $I=N_{G}(P,e)$ and $J=N_{H}(Q,f)$ and let $\hat{e}$ and $\hat{f}$ denote the unique block idempotents of $k[N_{G}(P)]$ and $k[N_{H}(Q)]$ which cover $e$ and $f$, respectively. Then,
@@
is a splendid Rickard complex for blocks algebras $k[N_{G}(P)]\hat{e}$ and $k[N_{H}(Q)]\hat{f}$.
@@
Given any Brauer pair $(P,e)$ of $kG$, we have a Morita equivalence ${}_{k[N_{G}(P)]\hat{e}}\mathbf{mod}\to{}_{kIe}\mathbf{mod}$, where $\hat{e}$ is the unique block of $kN_{G}(P)$ covering $e$. Explicitly, $\hat{e}=\operatorname{tr}^{N_{G}(P)}_{I}(e)$, and the Morita equivalence is induced by the $(kIe,k[N_{G}(P)]\hat{e})$-bimodule $e[N_{G}(P)]\hat{e}$ (see [15]). The result now follows since
@@
9. Structural properties of splendid Rickard complexes
@@
[4, Theorem 14.1, 14.3] Let $\gamma$ be a $p$-permutation equivalence between blocks $A$ and $B$ of $kG$ and $kH$, respectively. Let $N$ be an indecomposable $(A,B)$-bimodule appearing in $\gamma$. Then $\mathcal{B}\mathcal{P}(N)\subseteq\mathcal{B}\mathcal{P}(\gamma)$. Moreover, there is a unique indecomposable $(A,B)$-bimodule $M$ appearing in $\gamma$ with multiplicity ±$\pm 1$ whose vertex is of the form $\Delta(D,\phi,E)$ where $D$ is a defect group of $A$. If $M\not\cong N$, then $\mathcal{B}\mathcal{P}(N)\subset\mathcal{B}\mathcal{P}(M)$.
The module $M$ as above is called the maximal module of $\gamma$. We begin proving the analogous statement with a block-wise reformulation of a theorem in [14].
We first state an oft-used lemma. The proof of the following folklore lemma is from correspondence from P. Webb.
@@
Let $M$ and $N$ be $p$-permutation $kG$-modules and let $f:M\to N$ be a $kG$-module homomorphism. Assume that $N=N_{1}\oplus\cdots\oplus N_{k}$ is a decomposition into indecomposable summands. For each $1\leq i\leq k$, let $(P_{i},e_{i})$ be a maximal $N_{i}$-Brauer pair. Then $f$ is split surjective if and only if $e_{i}\cdot f(P_{i})$ is surjective for all $1\leq i\leq k$.
@@
The forward implication is immediate. We induct on the number of indecomposable summands of $N$. For the base case, suppose that $N$ is indecomposable. Suppose that $(P,e)$ is a maximal $N$-Brauer pair and that $e\cdot f(P)$ is surjective. By Proposition 7.2, $P$ is a vertex of $N$ and therefore $N(P)$ is indecomposable. Since $(P,e)$ is an $N$-Brauer pair, $eN(P)\neq 0$ which implies that $eN(P)=N(P)$. Since $e\cdot f(P)$ is surjective, it follows that $f(P)$ is surjective as well. By [13, 5.10.7], $f$ is split surjective.
@@
Let $C$ be a chain complex of $p$-permutation $kG$-modules, and let $\mathcal{X}$ be a subset of $\mathcal{B}\mathcal{P}(kG)$ which is $G$-stable and closed (under $\leq$). The following are equivalent:
@@
(b) implies (a) is straightforward. Let $n$ be the minimum integer for which $C_{n}\neq 0$. For each nonzero $C_{n}$, write $C_{n}=X_{n}\oplus Y_{n}$, where $X_{n}$ consists of all indecomposable summands which have a maximal pair contained in $\mathcal{X}$. Thus, $Y_{n}$ consists of all indecomposable summands of $C_{n}$ which have a maximal Brauer pair which is not contained in $\mathcal{X}$. We construct a chain complex $D$ identical to $C$, except in degree $n+1$, where we set $D_{n+1}=X_{n}\oplus C_{n+1}$, and add the identity map on $X_{n}$ to the differential of $D_{n+1}$, $d_{n+1}$. Notice that $f\cdot X_{n}(Q)=0$ for all $(Q,f)\not\in\mathcal{X}$ since $\mathcal{X}$ is a $G$-stable ideal. Therefore, $f\cdot C(Q)\cong f\cdot D(Q)$. But since $f\cdot C(Q)$ is acyclic for all $(Q,f)\not\in\mathcal{X}$ by assumption, so is $f\cdot D(Q)$. Therefore $d_{n+1}$ composed with projection onto $Y_{n}$ is split surjective by Lemma 9.3, and by construction, it follows that $d_{n+1}$ is split surjective. Thus, we have that $D\cong D^{\prime}\oplus(C_{n}\xrightarrow{\sim}C_{n})$, where $D^{\prime}$ has lowest nonzero degree $n+1$. Note that $C$ is homotopy equivalent to a shift of the mapping cone of the chain complex homomorphism $D^{\prime}\to X_{n}$. Moreover, if $(Q,f)\not\in\mathcal{X}$, $f\cdot C(Q)\cong f\cdot D(Q)\simeq f\cdot D^{\prime}(Q)$, and these complexes are by assumption acyclic. We now perform an inductive argument as follows. If the length of $C$ is one, then (b) holds, since there is only one nonzero term which must vanish after applying the Brauer construction at all $(Q,f)\not\in\mathcal{X}$, and the rest follows by the previous lemma.
@@
Let $C$ be an indecomposable splendid Rickard complex for block algebras $A\subseteq kG$ and $B\subseteq kH$. Every trivial source bimodule occurring in $C$ has a maximal Brauer pair contained in $\mathcal{B}\mathcal{P}(C)$. In particular, the set consisting of all Brauer pairs for every trivial source bimodule appearing in $C$ and the set of $C$-Brauer pairs coincide.
@@
By Corollary 8.8, $\mathcal{B}\mathcal{P}(C)$ is a $G\times H$-stable ideal in $\mathcal{B}\mathcal{P}(k[G\times H])$. Thus, by the preceding theorem, every trivial source bimodule occuring in $C$ has a maximal Brauer pair contained in $\mathcal{B}\mathcal{P}(C)$. Denoting the set of all Brauer pairs which occur in any trivial source bimodule in $C$ by $\mathcal{X}$, it follows that $\mathcal{X}\subseteq\mathcal{B}\mathcal{P}(C)$. On the other hand, if $(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $C$-Brauer pair, then $eC(\Delta(P,\phi,Q))f\neq 0$. Thus, there exists a trivial source bimodule $M$ appearing in $C$ such that $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{B}\mathcal{P}(M)$. Thus, $\mathcal{B}\mathcal{P}(C)\subseteq\mathcal{X}$. ∎
@@
Acknowledgements: The authors would like to thank their advisor, Robert Boltje, for the many hours of guidance and support which made this paper possible. They would also like to thank Peter Webb for communicating and providing a proof of Lemma 9.2.

### v2

@@ 8. Brauer pairs for splendid Rickard equivalences
where for $\nu=(\sigma,(S,a),\tau))$, we set $X(\nu):=N_{I\times H}(\Delta(P,\sigma,S),e\otimes a^{*})$, $Y(\nu):=N_{H\times K}(\Delta(S,\tau,R),a\otimes d^{*})$, $C(\nu):=eC(\Delta(P,\sigma,S))a$ and $C^{*}(\nu):=aC^{*}(\Delta(S,\tau,R))d$. In particular, when $\nu=(\phi,(Q,f),c_{t}\circ\psi^{-1})$, we have $X(\nu)=X$, $Y(\nu)={}^{(t,1)}Y$, $C(\nu)=C(\lambda)$ and, by Proposition 3.4, $)C^{*}(\nu)\cong{}^{(t,1)}(C(\omega$$^{*})$. Thus, since $(\phi,(Q,f),c_{t}\circ\psi^{-1})\in\hat{\Omega}$, we conclude that
[…]
@@
As a corollary of the preceding theorem, we obtain that the Brauer pairs for a splendid Rickard equivalence and the Brauer pairs for its Lefschetz invariant coincide. This allows us to reinterpret the results of [BoPe20] to the setting of splendid equivalences.
@@
c.f. [BoPe20, Theorems 10.11, 11.2] Let $C$ be a splendid Rickard equivalence between $A$ and $B$ and let $\gamma:=\Lambda(C)$ be its corresponding $p$-permutation equivalence. Then $\mathcal{BP}(C)=\mathcal{BP}(\gamma)$. In particular,
@@
By Proposition 7.7 and Theorem 8.4, we may assume that $A$ and $B$ are blocks. Then the fact that $\mathcal{BP}(C)\subseteq\mathcal{BP}(\gamma)$ follows from Theorem 8.7. The other inclusion $\mathcal{BP}(\gamma)\subseteq\mathcal{BP}(C)$ was already established at the beginning of this section. Claims (a)-(c) above follow immediately from [BoPe20, Theorem 10.11]. Claim (d) follows from [BoPe20, Theorem 11.2]. ∎
Another easy corollary we obtain is that splendid Rickard equivalences induce local normalizer splendid equivalences, again an analogue of a result of Boltje–Perepelitsky.
@@
c.f. [BoPe20, Theorem 11.4] Let $C$ be a splendid Rickard equivalence between blocks $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(P,\phi,Q),e\otimes f^{*})$ be a $C$-Brauer pair. Set $I=N_{G}(P,e)$ and $J=N_{H}(Q,f)$ and let $\hat{e}$ and $\hat{f}$ denote the unique block idempotents of $k[N_{G}(P)]$ and $k[N_{H}(Q)]$ which cover $e$ and $f$, respectively. Then,
@@
is a splendid Rickard equivalence for blocks algebras $k[N_{G}(P)]\hat{e}$ and $k[N_{H}(Q)]\hat{f}$.
@@
[…]
@@
9. Structural properties of splendid Rickard equivalences
@@
[BoPe20, Theorem 14.1, 14.3] Let $\gamma$ be a $p$-permutation equivalence between blocks $A$ and $B$ of $kG$ and $kH$, respectively. Let $N$ be an indecomposable $(A,B)$-bimodule appearing in $\gamma$. Then $\mathcal{B}\mathcal{P}(N)\subseteq\mathcal{B}\mathcal{P}(\gamma)$. Moreover, there is a unique indecomposable $(A,B)$-bimodule $M$ appearing in $\gamma$ with multiplicity $\pm 1$ whose vertex is of the form $\Delta(D,\phi,E)$ where $D$ is a defect group of $A$. If $M\not\cong N$, then $\mathcal{B}\mathcal{P}(N)\subset\mathcal{B}\mathcal{P}(M)$.
The module $M$ as above is called the maximal module of $\gamma$. We begin proving the analogous statement with a block-wise reformulation of a theorem in [SM23].
We first state an oft-used lemma. The proof of the following folklore lemma is from correspondence from Peter Webb.
@@
Let $M$ and $N$ be $p$-permutation kGkG-modules and let $f:M\to N$ be a $kG$-module homomorphism. Assume that $N=N_{1}\oplus\cdots\oplus N_{k}$ is a decomposition into indecomposable summands. For each $1\leq i\leq k$, let $(P_{i},e_{i})$ be a maximal $N_{i}$-Brauer pair. Then $f$ is split surjective if and only if $e_{i}\cdot f(P_{i})$ is surjective for all $1\leq i\leq k$.
@@
The forward implication is immediate. We induct on the number of indecomposable summands of $N$. For the base case, suppose that $N$ is indecomposable. Suppose that $(P,e)$ is a maximal $N$-Brauer pair and that $e\cdot f(P)$ is surjective. By Proposition 7.2, $P$ is a vertex of $N$ and therefore $N(P)$ is indecomposable. Since $(P,e)$ is an $N$-Brauer pair, $eN(P)\neq 0$ which implies that $eN(P)=N(P)$. Since $e\cdot f(P)$ is surjective, it follows that $f(P)$ is surjective as well. By [Li181, Proposition 5.10.7], $f$ is split surjective; we may use this proposition by passing to duals, as the internal hom $\operatorname{Hom}_{k}(-,k)$ is an antiequivalence of $K^{b}({}_{kG}\mathbf{mod})$.
@@
Let $C$ be a chain complex of $p$-permutation kGkG-modules, and let $\mathcal{X}$ be a subset of $\mathcal{B}\mathcal{P}(kG)$ which is $G$-stable and closed (under $\leq$). The following are equivalent:
@@
(b) implies (a) is straightforward. Let $n$ be the minimum integer for which $C_{n}\neq 0$. For each nonzero $C_{n}$, write $C_{n}=X_{n}\oplus Y_{n}$, where $X_{n}$ consists of all indecomposable summands which have a maximal pair contained in $\mathcal{X}$. Thus, $Y_{n}$ consists of all indecomposable summands of $C_{n}$ which have a maximal Brauer pair which is not contained in $\mathcal{X}$. We construct a chain complex $D$ identical to $C$, except in degree $n+1$, where we set $D_{n+1}=X_{n}\oplus C_{n+1}$, and add the identity map on $X_{n}$ to the differential of $D_{n+1}$, $d_{n+1}$. Notice that $f\cdot X_{n}(Q)=0$ for all $(Q,f)\not\in\mathcal{X}$ since $\mathcal{X}$ is a $G$-stable ideal. Therefore, $f\cdot C(Q)\cong f\cdot D(Q)$. But since $f\cdot C(Q)$ is acyclic for all $(Q,f)\not\in\mathcal{X}$ by assumption, so is $f\cdot D(Q)$. Therefore $d_{n+1}$ composed with projection onto $Y_{n}$ is split surjective by Lemma 9.3, and by construction, it follows that $d_{n+1}$ is split surjective. Thus, we have that $D\cong D^{\prime}\oplus(C_{n}\xrightarrow{\sim}C_{n})$, where $D^{\prime}$ has lowest nonzero degree $n+1$.
Note that $C$ is homotopy equivalent to a shift of the mapping cone of the chain complex homomorphism $D^{\prime}\to X_{n}$. Moreover, if $(Q,f)\not\in\mathcal{X}$, $f\cdot C(Q)\cong f\cdot D(Q)\simeq f\cdot D^{\prime}(Q)$, and these complexes are by assumption acyclic. We now perform an inductive argument as follows. If the length of $C$ is one, then (b) holds, since there is only one nonzero term which must vanish after applying the Brauer construction at all $(Q,f)\not\in\mathcal{X}$, and the rest follows by the previous lemma.
@@
Let $C$ be an indecomposable splendid Rickard equivalence for block algebras $A\subseteq kG$ and $B\subseteq kH$ and let $(\Delta(D,\phi,E),e\otimes f^{*})$ be a maximal $C$-Brauer pair. Every trivial source bimodule occurring in $C$ has a maximal Brauer pair contained in $(\Delta(D,\phi,E),e\otimes f^{*})$. In particular, the set consisting of all Brauer pairs for every trivial source bimodule appearing in $C$ and the set of $C$-Brauer pairs coincide.
@@
By Corollary 8.8, $\mathcal{B}\mathcal{P}(C)$ is a $G\times H$-stable ideal in $\mathcal{B}\mathcal{P}(k[G\times H])$ and any two maximal $C$-Brauer pairs are $G\times H$-conjugate. Thus, by the preceding theorem, every trivial source bimodule occurring in $C$ has a maximal Brauer pair contained in $(\Delta(D,\phi,E),e\otimes f^{*})$. Denoting the set of all Brauer pairs which occur in any trivial source bimodule in $C$ by $\mathcal{X}$, it follows that $\mathcal{X}\subseteq\mathcal{B}\mathcal{P}(C)$. On the other hand, if $(\Delta(P,\phi,Q),e\otimes f^{*})$ is a $C$-Brauer pair, then $eC(\Delta(P,\phi,Q))f\neq 0$. Thus, there exists a trivial source bimodule $M$ appearing in $C$ such that $(\Delta(P,\phi,Q),e\otimes f^{*})\in\mathcal{B}\mathcal{P}(M)$. Thus, $\mathcal{B}\mathcal{P}(C)\subseteq\mathcal{X}$. ∎
@@
Acknowledgements: The authors would like to thank their advisor, Robert Boltje, for the many hours of guidance and support which made this paper possible, and would like to thank Peter Webb for communicating Lemma 9.2.